# Gravitational lensing and shadows in the toron solution of Einstein’s equations using ray tracing methods

### v1

@@ 5.2. Light rings
From the previous examples, we observe that one of the effects of the NUT parameter $\ell$ is the lifting of the light rays from the equatorial plane. We expect this to have a consequence on the observed light rings with respect to the ones in Schwarzschild spacetime. Let us recall that in that case, the trajectory of the light rays on the photon sphere move along the sphere’s great circles. Thus the trajectory can be described as the intersection of the photon sphere and a plane passing through the origin. The first effect of a non-vanishing $\ell$ on light rings is that the size of the photon sphere is bigger, as shown by the formula (37) and Fig. (17). The second effect is that as $\ell$ is increased, the light rings originally on the equatorial plane is lifted out of the plane as seen in Fig. 16; in fact, as observed in Fig. (17), these rings are small circles of the photon sphere.
@@
The light rays represented in Fig. (16) correspond to different spacetimes with a fixed $m=1$ and $\ell\in\{0,0.25,0.5,0.75,1\}$, but they are plotted together in order to visualize the influence of the gravimagnetic mass $\ell$. It can be observed that the trajectories of these light rays are still rings, but in this case they are small circles of the photon spheres, meaning that they are described by the intersection of a sphere and a plane not passing through the origin as shown in Figs. 16 and 17. Additionally, Fig. 17 shows the comparison of light rings in a Schwarzschild and a NUT spacetime, given analogous initial conditions. These results agree with the estimation of [15] for the deviation with respect to the equatorial plane (denoted therein as conicity), which is proportional to the NUT parameter $\ell$.
@@ 5.3. Shadows and apparent image of the background
It has been shown in [14] that the shadows of NUT black holes are circular. Given a fixed $m$, the size of the shadow increases in dependence of $\ell$, as inferred from formula (37). With the help of ray tracing techniques, one can additionally study the effects of $\ell$ on the apparent image of the background space. For example, we can place an observer whose line of sight is directed towards the black hole, as shown by the diagram in Fig. 9.
@@
The diagram shown in Fig. 19 corresponds to the Schwarzschild case, but the same phenomenon is observed in Kerr spacetimes as well. The only difference is that the shadow is no longer circular, and it shifts to the right of the screen. Subsequent rings would be observed if the image is magnified near the shadow. The first Einstein ring acts as a border between the primary and secondary image of the background space. However, although there is a border between the primary and secondary images in NUT spacetimes, they do not originate from the same source. Moreover, since the light rays are either trapped or repelled by the NUT black hole, as shown in Figs. 14 and 15, the region behind the black hole is a blind spot, meaning that the light of the objects right behind the black hole will not reach the observer (at least in the primary and secondary images). As the gravimagnetic mass $\ell$ increases, the size of the hidden region will also increase. In contrast, the celestial sphere is visible in its entirety in Schwarzschild and Kerr spacetimes, although the image appears deformed the closer they are to the Einstein ring. The effect of the gravimagnetic mass on the apparent image of the background space is that this image will be twisted as $\ell$ increases, analogous to the observations in the spacetimes studied in [24], in which the metric component $g_{t\phi}$ is skew-symmetric with respect to the equatorial plane, while the remaining components are even. To highlight the purely gravimagnetic effects, we show the extreme case with $m=0$ and $\ell\neq 0$ in Fig. 19.
@@
In order to visualize the appearance of the phenomenon discussed above as the parameter $\ell$ increases in a more quantitative manner, we assign an artificial coloring to the celestial sphere following the colors of the X-ray map of the sky in Fig. 11 obtained by the eROSITA telescope. The line of sight of the observer is pointing towards the center of the galaxy (center of the X-ray map), which would correspond to $(\phi,\theta)=(\pi,0)$ in the celestial sphere coordinates shown in 12, but with a black hole in between.
@@
To explain why a region of the celestial sphere is hidden when $\ell\neq 0$, we show the backwards tracing of some light rays reaching the observer in both Schwarzschild and NUT spacetimes in Fig. 21. We plot the light rays reaching the screen with a constant angular aperture, i.e., those pixels drawing a perfect circle on the screen of the virtual camera. In Schwarzschild, each of these light rays travels on a plane and will eventually cross again, as observed in the plot. This behaviour gives rise to the Einstein ring. However, if $\ell\neq 0$ we observe that these light rays are twisted about the line of sight (the $x$-axis in this example) and they will never cross again, unlike those in Schwarzschild’s spacetime. This is the reason for the existence of a blind region behind the NUT black hole in the primary and secondary images. We color one of the light rays in green in order to observe the twist of a single light ray about the $x$-axis. The light rays shown in the figure would correspond to the pixels showing the secondary image of the background space.
@@ 6. Geodesics in the toron spacetime
In this section we present simulations analogous to those presented in the preceding section, both for individual geodesics and for the apparent images of extended objects. First, we determine the solutions of the IVP (34) with the same initial conditions used for Figs. 14 and 15 in order to illustrate the similarities and differences between toron and NUT spacetimes, especially the case with $m=0$ and $\ell\neq 0$, since the disk described be the toron solution can be interpreted as a massless object, as discussed before.
@@ 6.1. Initially-parallel light-rays
The value of $dp^{z}/ds|_{s=0}$ for a given initial condition will indicate whether the light ray will stay on the plane (if this value vanishes) or it will move in upwards (resp. downwards) direction if this value is positive (resp. negative). Let us recall that for initial conditions of this type, $dp^{z}/ds|_{s=0}$ vanishes identically in Kerr spacetimes, while in NUT spacetimes it only depends on the sign of the initial $p^{\phi}_{0}$, as indicated by formula (35). However, such a formula does not hold in toron spacetimes since the metric functions are not symmetric in $z$ and one would need to numerically compute the full expression for $dp^{z}/ds|_{s=0}$ to estimate the deviation. Therefore, numerical computations as shown in Fig. 23 are necessary not only because of this difficulty, but also because of the unique behavior of the toron metric functions in the vicinity of the gravitating object.
@@
The effect of the angular momentum in addition to the gravimagnetic mass is further observed in Fig. 24. Although the formula (36) holds in toron spacetimes, determining the deviation of the light rays based solely on $z$ is not possible; one would need to numerically evaluate $\partial_{\rho}A$ at the initial position. The deviation of these light rays can be observed in Fig. 24. The dragging direction in toron spacetimes is clockwise for some light rays and counterclockwise for others, analogous to NUT spacetimes. The difference is that the dragging effect is stronger when $z_{0}<0$, which is explained by the fact that the magnitude of $A$ is bigger when $z$ is negative, which can be observed in Fig. 5.
Since the disk is a source of angular momentum contributing to the dragging effect in counterclockwise direction, compared to spacetimes with only gravimagnetic mass, the deviation of a light ray that would otherwise be dragged in clockwise direction is slowed down and it is accelerated if it were to be dragged in clockwise direction. This is well observed by comparing the $xy$ projections in Figs. 15 and 24. As expected for toron spacetimes, the symmetry of the $xy$ projection of the geodesics with respect to the $x$-axis is lost and an additional dragging component in counterclockwise direction is observed.
@@ 6.2. Search for fundamental photon orbits
Here we are trying to find fundamental photon orbits in toron spacetime with $\alpha=0.3$. As in Kerr-NUT spacetimes, the fundamental photon orbits with maximal distance from the object correspond to a retrograde photon passing the equatorial plane with zero $z$-component of velocity. This orbit is shown on the left-hand side of Fig 25. However, unlike Kerr-NUT spacetimes, it is not possible to decouple the geodesic equations and thus, determining the shape of the fundamental orbit is a less trivial task. Moreover, light rays have unstable orbits, meaning that tiny perturbations in the initial conditions of the IVP (34) give rise to quite different outcomes at large times, as seen in Fig. 25, which shows two particular photon orbits being approached by three photons whose initial conditions differ only slightly. In order to see how some photons are eventually trapped by the disk, the bottom part of Fig. 25 shows the same trajectories in a diagram with the ergosphere. The size of the fundamental photon orbit depends on the magnitude of the parameter $\alpha$ of a given toron spacetime.
@@ 6.3. Apparent image of a disk
We continue the discussion by showing simulations analogous to the images in Fig. 20, in which a black hole is in front of the observer and therefore, it casts a shadow. To do this, we assume that the disk is not emitting light in order to observe a full shadow as in black hole spacetimes. Since there is an angular momentum component, the shadow is not circular in toron spacetimes. Qualitatively, the effect of the $\alpha$ parameter on the apparent image of the celestial sphere is similar to the effect of the gravimagnetic mass $\ell$ in NUT spacetimes, as one can observe by comparing Figs. 20 and 29; in particular, the case with $m=0$ and $\ell\neq 0$. The main consequence is the twisting of the apparent image of the background space, analogous to the observed phenomenon in NUT spacetimes and the spacetimes studied in [24]. However, there are differences as well, since the disk is also a source of angular momentum as discussed above. This is mainly observed in the shifting of the shadow to the right. Nevertheless, the dominant effect is that of the gravimagnetic mass component, since the primary and secondary images of the upper and lower halves of the celestial sphere show more resemblance to the image shown in Fig. 19 corresponding to the NUT spacetime with $m=0$ and $\ell\neq 0$. To obtain these simulations, we use the same values chosen for Fig. 20, which would produce angles of view of approximately $\delta_{H}=\delta_{V}=120^{\circ}$ in Minkowski spacetime. Namely, the observer is placed at a distance $R_{c}=20$ from the center of the disk, its line of sight has an inclination $\psi=90^{\circ}$ with respect to the symmetry axis and the radius of the celestial sphere is set to $R_{\infty}=100$; the screen has a width and height $d_{H}=d_{V}=0.1$, a resolution of $300\times 300$ pixels and a focal length $d_{L}=0.04$.
@@ 7. Conclusion
In this paper we have presented a discussion of physical properties of toron solutions to the stationary axisymmetric Einstein equations in vacuum. They have vanishing mass and a non-zero NUT parameter. Based on the numerical experiments, the free parameter $\alpha$ in the solution can be interpreted as a source of both gravimagnetic mass and angular momentum. One key difference with respect to well-known spacetimes is the existence of an infinite number of ergospheres in these spacetimes, all of which have a toroidal shape.
@@
The physical properties of these spacetimes are explored via ray tracing. Well-known results of ray tracing in Kerr and NUT spacetimes are recalled as a reference for similar effects in toron spacetimes. It was shown numerically that light rays initially moving in the equatorial plane are deflected in upwards or downwards directions similarly to NUT spacetimes. As in any stationary axisymmetric spacetime, light rays initially moving in a purely radial direction (i.e. with $p^{z}_{0}=p^{\phi}_{0}=0$) are dragged in a direction depending on the sign of the metric function $A$. The existence of fundamental photon orbits and corresponding photon surfaces is inferred numerically based on the observation of individual light rays almost orbiting on the surface of a sphere (then eventually falling into the gravitating object or escaping to infinity), as well as the observation of shadows similar to those appearing in black hole spacetimes.

### v2

@@ 5.2. Light rings
From the previous examples, we observe that one of the effects of the NUT parameter $\ell$ is the lifting of the light rays from the equatorial plane. We expect this to have a consequence on the observed light rings with respect to the ones in Schwarzschild spacetime. Let us recall that in that case, the trajectory of the light rays on the photon sphere move along the sphere’s great circles. Thus the trajectory can be described as the intersection of the photon sphere and a plane passing through the origin. The first effect of a non-vanishing $\ell$ on light rings is that the size of the photon sphere is bigger, as shown by the formula (4.3) and Fig. (15). The second effect is that as $\ell$ is increased, the light rings originally on the equatorial plane is lifted out of the plane as seen in Fig. 14; in fact, as observed in Fig. (15), these rings are small circles of the photon sphere.
@@
The light rays represented in Fig. (14) correspond to different spacetimes with a fixed $m=1$ and $\ell\in\{0,0.25,0.5,0.75,1\}$, but they are plotted together in order to visualize the influence of the gravimagnetic mass $\ell$. It can be observed that the trajectories of these light rays are still rings, but in this case they are small circles of the photon spheres, meaning that they are described by the intersection of a sphere and a plane not passing through the origin as shown in Figs. 14 and 15. Additionally, Fig. 15 shows the comparison of light rings in a Schwarzschild and a NUT spacetime, given analogous initial conditions. These results agree with the estimation of [15] for the deviation with respect to the equatorial plane (denoted therein as conicity), which is proportional to the NUT parameter $\ell$.
@@ 5.3. Shadows and apparent image of the background
It has been shown in [14] that the shadows of NUT black holes are circular. Given a fixed $m$, the size of the shadow increases in dependence of $\ell$, as inferred from formula (4.3). With the help of ray tracing techniques, one can additionally study the effects of $\ell$ on the apparent image of the background space. For example, we can place an observer whose line of sight is directed towards the black hole, as shown by the diagram in Fig. 7.
@@
The diagram shown in Fig. 17 corresponds to the Schwarzschild case, but the same phenomenon is observed in Kerr spacetimes as well. The only difference is that the shadow is no longer circular, and it shifts to the right of the screen. Subsequent rings would be observed if the image is magnified near the shadow. The first Einstein ring acts as a border between the primary and secondary image of the background space. However, although there is a border between the primary and secondary images in NUT spacetimes, they do not originate from the same source. Moreover, since the light rays are either trapped or repelled by the NUT black hole, as shown in Figs. 12 and 13, the region behind the black hole is a blind spot, meaning that the light of the objects right behind the black hole will not reach the observer (at least in the primary and secondary images). As the gravimagnetic mass $\ell$ increases, the size of the hidden region will also increase. In contrast, the celestial sphere is visible in its entirety in Schwarzschild and Kerr spacetimes, although the image appears deformed the closer they are to the Einstein ring. The effect of the gravimagnetic mass on the apparent image of the background space is that this image will be twisted as $\ell$ increases, analogous to the observations in the spacetimes studied in [24], in which the metric component $g_{t\phi}$ is skew-symmetric with respect to the equatorial plane, while the remaining components are even. To highlight the purely gravimagnetic effects, we show the extreme case with $m=0$ and $\ell\neq 0$ in Fig. 17.
@@
In order to visualize the appearance of the phenomenon discussed above as the parameter $\ell$ increases in a more quantitative manner, we assign an artificial coloring to the celestial sphere following the colors of the X-ray map of the sky in Fig. 9 obtained by the eROSITA telescope. The line of sight of the observer is pointing towards the center of the galaxy (center of the X-ray map), which would correspond to $(\phi,\theta)=(\pi,0)$ in the celestial sphere coordinates shown in 10, but with a black hole in between.
@@
To explain why a region of the celestial sphere is hidden when $\ell\neq 0$, we show the backwards tracing of some light rays reaching the observer in both Schwarzschild and NUT spacetimes in Fig. 19. We plot the light rays reaching the screen with a constant angular aperture, i.e., those pixels drawing a perfect circle on the screen of the virtual camera. In Schwarzschild, each of these light rays travels on a plane and will eventually cross again, as observed in the plot. This behavior gives rise to the Einstein ring. However, if $\ell\neq 0$ we observe that these light rays are twisted about the line of sight (the $x$-axis in this example) and they will never cross again, unlike those in Schwarzschild’s spacetime. This is the reason for the existence of a blind region behind the NUT black hole in the primary and secondary images. We color one of the light rays in green in order to observe the twist of a single light ray about the $x$-axis. The light rays shown in the figure would correspond to the pixels showing the secondary image of the background space.
@@ 6. Geodesics in the toron spacetime
In this section we present simulations analogous to those presented in the preceding section, both for individual geodesics and for the apparent images of extended objects. First, we determine the solutions of the IVP (B.5) with the same initial conditions used for Figs. 12 and 13 in order to illustrate the similarities and differences between toron and NUT spacetimes, especially the case with $m=0$ and $\ell\neq 0$, since the disk described be the toron solution can be interpreted as a massless object, as discussed before.
@@ 6.1. Initially-parallel light-rays
The value of $dp^{z}/ds|_{s=0}$ for a given initial condition will indicate whether the light ray will stay on the plane (if this value vanishes) or it will move in upwards (resp. downwards) direction if this value is positive (resp. negative). Let us recall that for initial conditions of this type, $dp^{z}/ds|_{s=0}$ vanishes identically in Kerr spacetimes, while in NUT spacetimes it only depends on the sign of the initial $p^{\phi}_{0}$, as indicated by formula (4.1). However, such a formula does not hold in toron spacetimes since the metric functions are not symmetric in $z$ and one would need to numerically compute the full expression for $dp^{z}/ds|_{s=0}$ to estimate the deviation. Therefore, numerical computations as shown in Fig. 21 are necessary not only because of this difficulty, but also because of the unique behavior of the toron metric functions in the vicinity of the gravitating object.
@@
The effect of the angular momentum in addition to the gravimagnetic mass is further observed in Fig. 22. Although the formula (4.2) holds in toron spacetimes, determining the deviation of the light rays based solely on $z$ is not possible; one would need to numerically evaluate $\partial_{\rho}A$ at the initial position. The deviation of these light rays can be observed in Fig. 22. The dragging direction in toron spacetimes is clockwise for some light rays and counterclockwise for others, analogous to NUT spacetimes. The difference is that the dragging effect is stronger when $z_{0}<0$, which is explained by the fact that the magnitude of $A$ is bigger when $z$ is negative, which can be observed in Fig. 5.
Since the disk is a source of angular momentum contributing to the dragging effect in counterclockwise direction, compared to spacetimes with only gravimagnetic mass, the deviation of a light ray that would otherwise be dragged in clockwise direction is slowed down and it is accelerated if it were to be dragged in clockwise direction. This is well observed by comparing the $xy$ projections in Figs. 13 and 22. As expected for toron spacetimes, the symmetry of the $xy$ projection of the geodesics with respect to the $x$-axis is lost and an additional dragging component in counterclockwise direction is observed.
@@ 6.2. Search for fundamental photon orbits
Here we are trying to find fundamental photon orbits in toron spacetime with $\alpha=0.3$. As in Kerr-NUT spacetimes, the fundamental photon orbits with maximal distance from the object correspond to a retrograde photon passing the equatorial plane with zero $z$-component of velocity. This orbit is shown on the left-hand side of Fig 23. However, unlike Kerr-NUT spacetimes, it is not possible to decouple the geodesic equations and thus, determining the shape of the fundamental orbit is a less trivial task. Moreover, light rays have unstable orbits, meaning that tiny perturbations in the initial conditions of the IVP (B.5) give rise to quite different outcomes at large times, as seen in Fig. 23, which shows two particular photon orbits being approached by three photons whose initial conditions differ only slightly. In order to see how some photons are eventually trapped by the disk, the bottom part of Fig. 23 shows the same trajectories in a diagram with the ergosphere. The size of the fundamental photon orbit depends on the magnitude of the parameter $\alpha$ of a given toron spacetime.
@@ 6.3. Apparent image of a disk
We continue the discussion by showing simulations analogous to the images in Fig. 18, in which a black hole is in front of the observer and therefore, it casts a shadow. To do this, we assume that the disk is not emitting light in order to observe a full shadow as in black hole spacetimes. Since there is an angular momentum component, the shadow is not circular in toron spacetimes. Qualitatively, the effect of the $\alpha$ parameter on the apparent image of the celestial sphere is similar to the effect of the gravimagnetic mass $\ell$ in NUT spacetimes, as one can observe by comparing Figs. 18 and 27; in particular, the case with $m=0$ and $\ell\neq 0$. The main consequence is the twisting of the apparent image of the background space, analogous to the observed phenomenon in NUT spacetimes and the spacetimes studied in [24]. However, there are differences as well, since the disk is also a source of angular momentum as discussed above. This is mainly observed in the shifting of the shadow to the right. Nevertheless, the dominant effect is that of the gravimagnetic mass component, since the primary and secondary images of the upper and lower halves of the celestial sphere show more resemblance to the image shown in Fig. 17 corresponding to the NUT spacetime with $m=0$ and $\ell\neq 0$. To obtain these simulations, we use the same values chosen for Fig. 18, which would produce angles of view of approximately $\delta_{H}=\delta_{V}=120^{\circ}$ in Minkowski spacetime. Namely, the observer is placed at a distance $R_{c}=20$ from the center of the disk, its line of sight has an inclination $\psi=90^{\circ}$ with respect to the symmetry axis and the radius of the celestial sphere is set to $R_{\infty}=100$; the screen has a width and height $d_{H}=d_{V}=0.1$, a resolution of $300\times 300$ pixels and a focal length $d_{L}=0.04$.
@@ 7. Conclusion
In this paper we have presented a discussion of physical properties of toron solutions to the stationary axisymmetric Einstein equations in vacuum. They have vanishing mass and a non-zero NUT parameter. The parameter $\alpha$ in the solution determines both gravimagnetic mass and angular momentum. One key difference with respect to well-known spacetimes is the existence of an infinite number of ergospheres in these spacetimes, all of which have a toroidal shape.
The physical properties of these spacetimes are explored via ray tracing. Well-known results of ray tracing in Kerr and NUT spacetimes are recalled as a reference for similar effects in toron spacetimes. The main findings are as follows:
- Role of a NUT parameter: as a consequence of the presence of a NUT parameter the light rays initially moving in the equatorial plane deviate in upwards or downwards directions. This behavior is numerically observed in toron spacetimes similarly to NUT spacetimes. This is an indication that the so-called gravimagnetic mass is probably not relevant in known astrophysical situations since no light ray deviation of such type was observed so far.
- Frame dragging: as in any stationary axisymmetric spacetime, light rays initially moving in a purely radial direction (i.e. with $p^{z}_{0}=p^{\phi}_{0}=0$) are dragged in a direction depending on the sign of the metric function $A$.
@@
- Fundamental photon orbits: the existence of fundamental photon orbits and corresponding photon surfaces is inferred numerically based on the observation of individual light rays almost orbiting on some surface of a spherical topology (then eventually falling into the gravitating object or escaping to infinity), as well as the observation of shadows similar to those appearing in black hole spacetimes. Due to a lack of symmetry in toron spacetimes compared to NUT spacetimes, it is more straightforward to identify them in the latter.